OpenStudy (anonymous):

d/dx(ln[e^(2(x)^1/2)cot^2((x)^1/2))

OpenStudy (anonymous):

My beautiful eyes!

OpenStudy (anonymous):

lol,what?

OpenStudy (anonymous):

\[\frac{d}{dx}\left(\ln\left(e^{2x^{\frac{1}{2}}}\cot^2\left(x^{\frac{1}{2}}\right)\right)\right)\]This?

OpenStudy (anonymous):



OpenStudy (anonymous):

the others are right

OpenStudy (campbell_st):

simplify the problem using log laws \[\ln (e^{2\sqrt{x}}) + \ln (\cot^2(\sqrt{x}))\]

OpenStudy (campbell_st):

the 1st term becomes \[2x^{1/2}\]

OpenStudy (anonymous):

yeah did that

OpenStudy (anonymous):

why 2x^1/2? that's where am stock?

OpenStudy (campbell_st):

so the derivative is \[1/\sqrt{x}\]

OpenStudy (anonymous):

ok

OpenStudy (campbell_st):

the log of the exponential is just the power

OpenStudy (campbell_st):

you just need to find the derivative of the cot^2 stuff

OpenStudy (anonymous):

yeah am getting this